OpenStudy (anonymous):

x = 3 pi over 4, x = 5 pi over 4 x = pi over 2, x = 3 pi over 2 x = 3 pi over 4, x = 5 pi over 4, x = 7 pi over 4 x = pi over 2, x = 3 pi over 2, x = 5 pi over 2

OpenStudy (xapproachesinfinity):

So what's your question?

OpenStudy (anonymous):

\[X = 3 x \frac{ \Pi}{ 4 }, X = 5 x \frac{ n }{ 4 }\] \[X = \frac{ 3\Pi }{ 4}, x= \frac{ 5\Pi }{ 4 }\]

OpenStudy (anonymous):

Your answer is at the bottom (Hope this helped)

OpenStudy (xapproachesinfinity):

there is no question at all to begin with!!! pls post question properly. many guys here post none sense! and solve with not info at all